OpenStudy (anonymous):

r-s=-3 r+s=9 solve the equation (messed up question last time)

OpenStudy (anonymous):

substitution solve the first equation in terms of r or s then substitute it into the 2nd equation and solve for teh remaining variable then substitute the variable you solved back into the original equations to solve for the other variable

OpenStudy (anonymous):

????????

OpenStudy (anonymous):

first equation r-s=-3 we solve it in terms of s r=s-3 we substitute for r in the 2nd equation s-3+s=9 we solve for s

OpenStudy (anonymous):

oh okay got it now thanks!

OpenStudy (anonymous):

once you have s substitute that back into the original equation r-s=-3 or r+s=9 then solve for r